An apple orchard contains 96 trees. The number of trees in each row is 4 less than twice the number of rows. Find the number of rows and the number of trees per row
Answer by ewatrrr(24785)

n(n-4)= 96
n^2- 4n = 96
(n+8)(n-12) = 0, n = 12 (positive solution)
12 rows, 8 trees per row